Find a direct relationship between x and y
x = t - 5 and y = t^2 - 6t

Answer by ikleyn(52847)

.
Express  t = x + 5  and substitute it into the quadratic expression for  t


   y = t^2 - 6t = (x+5)^2 - 6*(x+5).


Complete it and transform into another form on your own, if you need it.